# A high-order recombination algorithm for weak approximation of stochastic differential equations

### v1

@@ 1.3.2. High-order recombination measures and their construction procedures
$\begin{split}\mu^{(\rm{Alg},\Delta,\rm{Rec})}_{t_{0}}&=\delta_{x_{0}}\\ \nu^{(i)}&=\hat{Q}_{(s_{i})}^{(\rm{Alg},\Delta)}\left(\mu_{t_{i-1}}^{(\rm{Alg},\Delta,\rm{Rec})}\right)\\ \mu^{(\rm{Alg},\Delta,\rm{Rec})}_{t_{i}}&=\sum_{k=1}^{l_{i}}\nu^{(i)}\left(U^{(i)}_{k}\right)\tilde{\nu}^{(i)}_{k},\end{split}$ (11)
where $\left\{U^{(i)}_{k}\right\}_{k=1}^{l_{i}}$ is a set of subsets of $\operatorname{supp}\left(\nu^{(i)}\right)$ that satisfies
@@
$\tilde{\nu}^{(i)}_{k}\in\operatorname{red}_{G}\left(\left.\nu^{(i)}\right|_{U^{(i)}_{k}}\right)\quad\text{for $1\leq k\leq l_{i}$}$ (13)
and
@@
which is the induced probability measure on $U^{(i)}_{k}$ from $\nu^{(i)}=\sum_{j=1}^{m_{i}}\hat{w}_{j}^{(i)}\delta_{p^{(i)}_{j}}$. Henceforth, a set $\left\{U_{k}\right\}_{k=1}^{l}$ that satisfies
@@
is referred to as a patch division of $\operatorname{supp}(\nu^{(i)})$. Note that this procedure is interwoven with the tree-based simulation procedure and is performed simultaneously. Also note that the total weights assigned to each patch are invariant throughout the recombination procedure. We therefore refer to this as the weight of the patch $U_{k}^{(i)}$ and denote it by $\operatorname{wt}\left(U^{(i)}_{k}\right)$, that is
@@
In [17] they show that we can proceed with the high-order recombination while maintaining the order of discretisation by choosing appropriate patch radii
@@
We can find numerical examples of the high-order recombination method for practical problems only in [23]. In the paper they show numerically that even with ad hoc patch division, which is not proved to be free from the support explosion problem, the recombination method can still be effective for certain problems.
In this paper, we provide a detailed explanation of the algorithms corresponding to the aforementioned procedure. Specifically, the explicit construction of $\hat{Q}_{(s_{i})}^{(\rm{NV},\Delta)}$ is presented in Subsection 2.3, an algorithm for the patch division of $\operatorname{supp}(\nu^{(i)})$ into $\left\{U^{(i)}_{k}\right\}_{k=1}^{l_{i}}$ is proposed in Subsection 3.2, and the algorithm for finding $\tilde{\nu}^{(i)}_{k}\in\operatorname{red}_{G}\left(\left.\nu^{(i)}\right|_{U^{(i)}_{k}}\right)$, which is the same algorithm used in [23], is provided in Appendix 1.
@@ 1.4. Contribution of this paper
This paper presents a novel recursive patch division algorithm for high-order recombination that addresses the support explosion problem, making it practically feasible for real-world applications in mathematical finance. Our contributions can be summarized as follows
  • An improved version of the error estimation given in [17] is presented, which considers the patch weights $\operatorname{wt}\left(U^{(i)}_{k}\right)$.
  • A recursive algorithm is proposed, which divides the support set into patches, taking into account both the radii and the weights of the patches. The algorithm contains a control parameter that can be varied to change the degree of recombination. The recursive nature of this algorithm allows for efficient and adaptive patch division based on the refined error evaluation.
  • An example of the application of this algorithm to some real financial option pricing problems is presented. In this example, the algorithm successfully avoids the support explosion problem as expected. It also shows how changing the parameter controlling the degree of recombination works in this example. n
The remainder of this paper is organized as follows. In Section 2, we introduce the necessary notation and preliminary results on high-order discretisation schemes and high-order recombination. In Section 3, we present the refined error evaluation and the novel recursive patch division algorithm, which are the main contributions of this paper. In Section 4, we provide numerical results to demonstrate the performance of our algorithm.
@@ 2. Preliminary
In this section, we introduce some notations and previous studies on high-order discretisation schemes and the recombination algorithm.
@@ 2.1. Free Lie algebras
Let $A=\{0,1,\dots,d\}$ and define the set of all words consisting of elements of $A$ as $A^{+}=\bigcup_{k=1}^{\infty}A^{k}$ and also define the set $A^{\ast}=\{\emptyset\}\cup A^{+}$. $A^{\ast}$ becomes a semigroup with identity $\emptyset$ when endowed with the concatenating product $\cdot$’, that is for arbitrary elements $\left(a_{1},\dots,a_{m}\right)$ and $\left(b_{1},\dots,b_{n}\right)$ of $A^{\ast}$,
@@
To simplify the notation, we identify $(a_{1},a_{2},\ldots,a_{m})\in A^{\ast}$ with $a_{1}\cdot a_{2}\cdots a_{m}$. If there is no risk of confusion, $\cdot$’ is also omitted. That is, we would rather write $a_{1}a_{2}\ldots a_{m}$ instead of $(a_{1},a_{2},\ldots,a_{m})$ or $a_{1}\cdot a_{2}\cdots\cdot a_{m}$. For $\alpha=a_{1}a_{2}\dots a_{k}\in A^{\ast}$, we define the length of $\alpha$ as $\lvert\alpha\rvert=k$ and $\lvert\emptyset\rvert=0$. We also define the order of $\alpha$ as $\|\alpha\|=|\alpha|+\operatorname{card}\left(\{1\leq i\leq k\;|\;a_{i}=0\}\right)$. Let $A^{\ast}_{1}=A^{\ast}\backslash\{\emptyset,0\}$ and $A^{\ast}(m)=\{\alpha\in A^{\ast}\;|\;\|\alpha\|\leq m\}$ and $A^{\ast}_{1}(m)=\{\alpha\in A^{\ast}_{1}\;|\;\|\alpha\|\leq m\}$ respectively for $m\in\mathbb{N}$.
@@
Let $\mathbb{R}\langle A\rangle$ be the free $\mathbb{R}$-algebra with basis $A^{\ast}$ and $\mathbb{R}\langle\langle A\rangle\rangle$ be its completion, that is $\mathbb{R}\langle A\rangle=\left\{\left.\sum_{k=1}^{l}r_{k}\alpha_{k}\;\right|\;l\in\mathbb{N},r_{k}\in\mathbb{R},\alpha_{k}\in A^{\ast}\right\}$ and $\mathbb{R}\langle\langle A\rangle\rangle=\left\{\left.\sum_{\alpha\in A^{\ast}}r_{\alpha}\alpha\;\right|\;r_{\alpha}\in\mathbb{R}\right\}$. The latter is the set of all $\mathbb{R}$-coefficient formal series with basis $A^{\ast}$. For $R=\sum_{\alpha\in A^{\ast}}r_{\alpha}\alpha$ and $S=\sum_{\beta\in A^{\ast}}s_{\beta}\beta\in\mathbb{R}\langle A\rangle$, let their product $RS$ be defined by
$RS=\sum_{\theta=\alpha\beta\in A^{\ast}}r_{\alpha}s_{\beta}\theta,$
and their inner product $\langle R,S\rangle$ by
$\langle R,S\rangle=\sum_{\alpha=\beta}r_{\alpha}s_{\beta}.$
@@
We also define $\|R\|_{2}=\left(\langle R,R\rangle\right)^{1/2}$. The Lie bracket is defined to be $[R,S]=RS-SR$ as usual. ${\mathcal{L}}_{\mathbb{R}}(A)$ denotes the set of Lie polynomials in $\mathbb{R}\langle A\rangle$, that is the smallest $\mathbb{R}$-submodule of $\mathbb{R}\langle A\rangle$ including $A$ and closed under the Lie bracket.
@@
Let $\Phi$ be the homomorphism from $\mathbb{R}\langle A\rangle$ to the $\mathbb{R}$-algebra consisting of smooth differential operators over $\mathbb{R}^{N}$ such that
@@
$\begin{split}\Phi(\emptyset)&=\text{Id},\\ \Phi(\alpha)&=V_{a_{1}}V_{a_{2}}\cdots V_{a_{k}}\mbox{\hskip 5.69054pt for \hskip 5.69054pt}\alpha=a_{1}a_{2}\dots a_{k}\in A^{+}.\end{split}$ (18)
Here Id is the identity operator.
@@
We remark that $\Phi$ induces canonically a Lie algebra homomorphism from $\mathcal{L}_{\mathbb{R}}(A)$ to the Lie algebra of vector fields generated by $\left\{V_{0},\dots,V_{d}\right\}$. We denote $V_{R}=\Phi(R)$ for $R\in\mathbb{R}\langle A\rangle$ and $[\alpha]=[[\dots,[a_{1},a_{2}]\dots],a_{k}]$ for $\alpha=a_{1}a_{2}\dots a_{k}\in A^{\ast}$.
We remark that the following equality
$f(\exp(t\Phi(R))(x)=\Phi\left(j_{m}\left(\exp\left(tR\right)\right)\right)f(x)+O(t^{m+1})$ (19)
@@
holds for $R\in{\mathcal{L}}_{\mathbb{R}}(A)$ by virtue of the Taylor expansion formula.
@@ 2.2. High-order discretisation scheme
In this subsection, we introduce high-order discretisation schemes of SDEs and related results. $C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$ denotes the space of Lipschitz functions defined on $\mathbb{R}^{N}$, $\|\cdot\|_{\rm{Lip}}$ denotes the Lipschitz norm, that is
$\|f\|_{\rm{Lip}}=\sup_{x,y\in\mathbb{R}^{N},x\neq y}\frac{\left|f(x)-f(y)\right|}{\left|x-y\right|},$
and $\|\cdot\|_{\infty}$ denotes the uniform norm, that is $\left\|f\right\|_{\infty}=\sup_{x\in\mathbb{R}^{N}}\left|f(x)\right|$ for $f\in C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$.
@@
Let $X(t,x)$ be the $\mathbb{R}^{N}$-valued diffusion process defined by (1). For $f\in C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$ and $t\geq 0$, we define
$(P^{X}_{t}f)(x)=E\left[f(X(t,x))\right]$
and obtain a semigroup of linear operators $\{P^{X}_{t}\}_{t\geq 0}$ on $C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$.
At first, we present an approximation theorem for high-order discretisation scheme, based on [11, 12] and [13].
@@
Let $\ell$ be an integer. The vector fields $V_{0},V_{1},\dots,V_{d}$ are said to satisfy the UFG condition, if for all $\alpha\in A^{\ast}_{1}$, there exists $\left\{{\varphi}_{{\alpha},{\beta}}\>\left|\;\beta\in A^{\ast}_{1}\left(\ell\right)\right.\right\}\subset C^{\infty}_{b}(\mathbb{R}^{N};\mathbb{R})$ satisfying
@@
Hereinafter, $\ell$ is referred to as the degree of the UFG condition. Then, we define the condition on approximation operators as follows.
Theorem 1 (Smoothness of the semigroup $\left\{P^{X}_{t}\right\}_{t\geq 0}$ in certain directions [12]).
Assume that the vector fields $V_{0},V_{1},\dots,V_{d}$ satisfy the UFG condition and $X$ is the diffusion process defined as (1). Let $\alpha_{1},\alpha_{2},\dots,\alpha_{k}\in A^{\ast}_{1}$. Then, there exists a constant $C>0$ such that
$\left\|V_{[\alpha_{1}]}\cdots V_{[\alpha_{k^{\prime}}]}P^{X}_{t}V_{[\alpha_{k^{\prime}+1}]}\cdots V_{[\alpha_{k}]}f\right\|_{L^{p}(dx)}\leq Ct^{\left(\|\alpha_{1}\|+\cdots+\|\alpha_{k}\|\right)/2}\|f\|_{L^{p}(dx)}$
@@
Markov operator $\tilde{Q}_{(s)}$ is said to be an $m$-similar operator if there exist ${\mathcal{L}}_{\mathbb{R}}(A)$-valued random variables $L_{(s)}^{(1)},L_{(s)}^{(2)},\dots,L_{(s)}^{(K)}$ such that
@@
$E\left[\left|\left|j_{m}\left(L_{(s)}^{(k)}\right)\right|\right|_{2}^{n}\right]<\infty,\quad\left\langle L_{(s)}^{(k)},(0)\right\rangle=1\quad\text{for $n\in\mathbb{N}$, $k=1,2,\dots K$},$
@@
Here $B^{\circ\alpha}(s)$ is defined for $\alpha\in A^{*}$ as follows:
$B^{\circ\alpha}(t)=\begin{cases}1&\text{if $\alpha=\emptyset$},\\ \displaystyle{\int_{0}^{t}B^{\alpha^{\prime}}(s)\,\circ\mathrm{d}B^{a}(s)}&\text{if $\alpha=\alpha^{\prime}a$ where $\alpha^{\prime}\in A^{*}$ and $a\in A$}.\end{cases}$
The $m$-similar operator is intuitively understood as an approximation of $E[f\left(X\left(T,x\right)\right)]$ up to the $m$th order using a stochastic Taylor expansion of the solution of the SDE (1). We refer to $m$ as the degree of the discretisation method. Note that this definition has been slightly extended from the original definition used in [11, 12, 13].
Theorem 2 (Approximation theorem for high-order discretisation schemes [13]).
Assume that the vector fields of the SDE (1) satisfy the UFG condition, $X$ is the diffusion process defined by (1) and $Q_{(s)}$ is an $m$-similar operator. Then there exist positive constants $C$ and $C^{\prime}$ which depend only on $m$ and the vector fields $V_{0},V_{1},\dots,V_{d}$ and satisfy
$\left\|P^{X}_{s}f-Q_{(s)}f\right\|_{\infty}\leq Cs^{(m+1)/{2}}\left\|f\right\|_{\rm{Lip}}$
@@
$\left\|Q_{(s)}f\right\|_{\infty}\leq\exp(C^{\prime}s)\left\|f\right\|_{\infty}$
for all $f\in C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$.
Consequently, the total discretisation is bounded as follows. Let $T>0$ be a fixed time and consider a partition $\Delta$ of interval $[0,T]$ such that $\sharp\Delta=n$ as
$t_{0}=0<t_{1}<\dots<t_{n}=T.$
Then there exists a constant $C^{\prime\prime}$ depending only on $m$, the vector fields $V_{0},V_{1},\dots,V_{d}$, and $T$ such that
$\left\|P^{X}_{T}f-Q_{(s_{n})}\circ Q_{(s_{n-1})}\circ\cdots\circ Q_{(s_{1})}f\right\|_{\infty}\\ \leq C^{\prime\prime}\left\|f\right\|_{\rm{Lip}}\left(s_{n}^{1/2}+\sum_{j=m}^{m+1}\sum_{i=1}^{n-1}\frac{s_{i}^{(j+1)/2}}{(T-t_{i})^{j/2}}\right),$ (20)
@@
In the initial work of Kusuoka [11, 13], Theorem 2 is proved on the assumption of $f\in C_{b}^{\infty}(\mathbb{R}^{N};\mathbb{R})$. It is easily seen that this assumption can be relaxed to $f\in C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$, as pointed out in Section 3 of [18].

### v2

@@ 1.3.2. High-order recombination measures and their construction procedures
$\begin{split}\mu^{(\rm{Alg},\Delta,\rm{Rec})}_{t_{0}}&=\delta_{x_{0}},\\ \nu^{(i)}&=\hat{Q}_{(s_{i})}^{(\rm{Alg},\Delta)}\left(\mu_{t_{i-1}}^{(\rm{Alg},\Delta,\rm{Rec})}\right),\\ \mu^{(\rm{Alg},\Delta,\rm{Rec})}_{t_{i}}&=\sum_{k=1}^{l_{i}}\nu^{(i)}\left(U^{(i)}_{k}\right)\tilde{\nu}^{(i)}_{k},\end{split}$ (12)
where $\left\{U^{(i)}_{k}\right\}_{k=1}^{l_{i}}$ is a set of subsets of $\operatorname{supp}\left(\nu^{(i)}\right)$ satisfying
@@
and where
$\tilde{\nu}^{(i)}_{k}\in\operatorname{red}_{G}\left(\left.\nu^{(i)}\right|_{U^{(i)}_{k}}\right)\quad\text{for }1\leq k\leq l_{i},$ (14)
@@
which is the probability measure on $U^{(i)}_{k}$ induced by $\nu^{(i)}=\sum_{j=1}^{m_{i}}\hat{w}_{j}^{(i)}\delta_{p^{(i)}_{j}}$. Henceforth a set $\left\{U_{k}\right\}_{k=1}^{l}$ satisfying
@@
is referred to as a patch division of $\operatorname{supp}(\nu^{(i)})$. Note that this procedure is interwoven with the tree-based simulation and is carried out simultaneously. Also note that the total weight assigned to each patch remains invariant throughout the recombination procedure. We therefore refer to this as the weight of the patch $U_{k}^{(i)}$ denoted by $\operatorname{wt}\left(U^{(i)}_{k}\right)$. The following identity summarises the quantities introduced above:
@@
It is shown in [17] that high-order recombination can be performed in such a way that the order of discretisation is preserved by choosing appropriate patch radii
@@
Numerical examples of the high-order recombination method applied to practical problems appear only in [23]. It is shown numerically in that paper that, even when ad hoc patch division is employed—which has not been proved to avoid the support explosion problem—the recombination method can still be effective for certain problems.
This paper provides a detailed explanation of the algorithms corresponding to the aforementioned procedure. Specifically:
  • The explicit construction of $\hat{Q}_{(s_{i})}^{(\rm{NV},\Delta)}$ is presented in subsection 2.3.
  • An algorithm for the patch division of $\operatorname{supp}(\nu^{(i)})$ into $\left\{U^{(i)}_{k}\right\}_{k=1}^{l_{i}}$ is proposed in subsection 3.2.
  • The algorithm for finding $\tilde{\nu}^{(i)}_{k}\in\operatorname{red}_{G}\left(\left.\nu^{(i)}\right|_{U^{(i)}_{k}}\right)$—which is identical to the one used in [23]—is provided in Appendix 1.
@@ 1.4. Contribution of this paper
This paper presents a novel recursive patch division algorithm for high-order recombination that addresses the support explosion problem, thereby making it practically feasible for real-world applications in mathematical finance. The contributions of this paper are summarized as follows:
  • An improved version of the error estimation in [17] is presented, which incorporates the patch weights $\operatorname{wt}\left(U^{(i)}_{k}\right)$.
  • A recursive algorithm is proposed that divides the support set into patches, taking into account both the radii and the weights of the patches. The algorithm includes a control parameter that can be varied to change the degree of recombination. Its recursive structure enables efficient and adaptive patch division based on the refined error evaluation.
  • An example is provided demonstrating the application of the algorithm to realistic financial option pricing problems. In this example the algorithm successfully avoids the support explosion problem and illustrates how adjusting the control parameter affects the degree of recombination.
The remainder of this paper is organised as follows. Section 2 introduces the necessary notation and preliminary results on high-order discretisation schemes and high-order recombination. Section 3 presents the refined error evaluation and the novel recursive patch division algorithm, which constitute the main contributions of this paper. Section 4 provides numerical results that demonstrate the performance of the proposed algorithm.
@@ 2. Preliminary
This section introduces the notation and prior work on high-order discretisation schemes and the recombination algorithm.
@@ 2.1. Free Lie algebras
Let $A=\{0,1,\dots,d\}$ and define $A^{+}=\bigcup_{k=1}^{\infty}A^{k}$ as the set of all words consisting of elements of $A$, and define $A^{\ast}=\{\emptyset\}\cup A^{+}$. $A^{\ast}$ becomes a semigroup with identity $\emptyset$ when endowed with the concatenating product $\cdot$’, that is, for arbitrary elements $\left(a_{1},\dots,a_{m}\right)$ and $\left(b_{1},\dots,b_{n}\right)$ in $A^{\ast}$,
@@
To simplify the notation, we identify $(a_{1},a_{2},\ldots,a_{m})\in A^{\ast}$ with $a_{1}\cdot a_{2}\cdots a_{m}$. If there is no risk of confusion, we also omit the symbol $\cdot$’. That is, we prefer to write $a_{1}a_{2}\ldots a_{m}$ instead of $(a_{1},a_{2},\ldots,a_{m})$ or $a_{1}\cdot a_{2}\cdots\cdot a_{m}$. For $\alpha=a_{1}a_{2}\dots a_{k}\in A^{\ast}$, we define the length of $\alpha$ by $\lvert\alpha\rvert=k$, and set $\lvert\emptyset\rvert=0$. We also define the order of $\alpha$ as
$\|\alpha\|=\lvert\alpha\rvert+\operatorname{card}\left(\{1\leq i\leq k\;|\;a_{i}=0\}\right).$
Let $A^{\ast}_{1}=A^{\ast}\setminus\{\emptyset,0\}$ and define
$A^{\ast}(m)=\{\alpha\in A^{\ast}\;|\;\|\alpha\|\leq m\}\quad\text{and}\quad A^{\ast}_{1}(m)=\{\alpha\in A^{\ast}_{1}\;|\;\|\alpha\|\leq m\}$
for $m\in\mathbb{N}$.
@@
Let $\mathbb{R}\langle A\rangle$ be the free $\mathbb{R}$-algebra with basis $A^{\ast}$ and $\mathbb{R}\langle\langle A\rangle\rangle$ be its completion, that is
$\mathbb{R}\langle A\rangle=\left\{\left.\sum_{k=1}^{l}r_{k}\alpha_{k}\;\right|\;l\in\mathbb{N},r_{k}\in\mathbb{R},\alpha_{k}\in A^{\ast}\right\}$
and
$\mathbb{R}\langle\langle A\rangle\rangle=\left\{\left.\sum_{\alpha\in A^{\ast}}r_{\alpha}\alpha\;\right|\;r_{\alpha}\in\mathbb{R}\right\}.$
The latter is the set of all $\mathbb{R}$-coefficient formal series with basis $A^{\ast}$. The multiplication in $\mathbb{R}\langle\langle A\rangle\rangle$ is defined by extending the product in $\mathbb{R}\langle A\rangle$ continuously. More precisely, for such $R,S\in\mathbb{R}\langle\langle A\rangle\rangle$ written as
$R=\sum_{\alpha\in A^{\ast}}r_{\alpha}\alpha,\quad S=\sum_{\beta\in A^{\ast}}s_{\beta}\beta,$
their product is given by
$RS=\sum_{\gamma\in A^{\ast}}\left(\sum_{\begin{subarray}{c}\alpha,\beta\in A^{\ast}\\ \alpha\beta=\gamma\end{subarray}}r_{\alpha}s_{\beta}\right)\gamma.$
This product makes $\mathbb{R}\langle\langle A\rangle\rangle$ a unital non-commutative algebra over $\mathbb{R}$, in which $\mathbb{R}\langle A\rangle$ is a subalgebra. We also define $\|R\|_{2}=\left(\langle R,R\rangle\right)^{1/2}$. The Lie bracket is defined by $[R,S]=RS-SR$ as usual.
@@
${\mathcal{L}}_{\mathbb{R}}(A)$ denotes the set of Lie polynomials over $A$, that is, the smallest $\mathbb{R}$-submodule of $\mathbb{R}\langle A\rangle$ that contains $A$ and is closed under the Lie bracket [,][,][ , ].
@@
Let $\Phi$ be the homomorphism from $\mathbb{R}\langle A\rangle$ to the $\mathbb{R}$-algebra of smooth differential operators on $\mathbb{R}^{N}$, defined by
@@
$\begin{split}\Phi(\emptyset)&=\mathrm{Id},\\ \Phi(\alpha)&=V_{a_{1}}V_{a_{2}}\cdots V_{a_{k}}\quad\text{for $\alpha=a_{1}a_{2}\dots a_{k}\in A^{+}$.}\end{split}$ (19)
Here, $\mathrm{Id}$ denotes the identity operator.
@@
We remark that $\Phi$ canonically induces a Lie algebra homomorphism from $\mathcal{L}_{\mathbb{R}}(A)$ to the Lie algebra of vector fields generated by $\{V_{0},\dots,V_{d}\}$. We denote $V_{R}=\Phi(R)$ for $R\in\mathbb{R}\langle A\rangle$, and define $[\alpha]=[[[\dots[[a_{1},a_{2}],a_{3}],\dots],a_{k-1}],a_{k}]$ for $\alpha=a_{1}a_{2}\dots a_{k}\in A^{*}$, that is the right-nested Lie bracket of the letters in $\alpha$.
@@
We remark that the following approximation holds for $R\in{\mathcal{L}}_{\mathbb{R}}(A)$ by virtue of the Taylor expansion formula:
$f(\exp(t\Phi(R))(x))=\Phi\left(j_{m}\left(\exp(tR)\right)\right)f(x)+O(t^{m+1}).$ (20)
@@ 2.2. High-order discretisation scheme
In this subsection, we introduce high-order discretisation schemes for SDEs and related results.
Let $C_{\mathrm{Lip}}(\mathbb{R}^{N};\mathbb{R})$ denote the space of Lipschitz continuous functions on $\mathbb{R}^{N}$, and let $\|\cdot\|_{\mathrm{Lip}}$ denote the Lipschitz norm defined by
$\|f\|_{\mathrm{Lip}}=\sup_{x,y\in\mathbb{R}^{N},\,x\neq y}\frac{|f(x)-f(y)|}{|x-y|},$
and let $\|\cdot\|_{\infty}$ denote the uniform norm defined by
$\|f\|_{\infty}=\sup_{x\in\mathbb{R}^{N}}|f(x)|$
for $f\in C_{\mathrm{Lip}}(\mathbb{R}^{N};\mathbb{R})$.
@@
Let $X(t,x)$ be the $\mathbb{R}^{N}$-valued diffusion process defined by (1). For $f\in C_{\mathrm{Lip}}(\mathbb{R}^{N};\mathbb{R})$ and $t\geq 0$, define
$(P^{X}_{t}f)(x)=E[f(X(t,x))]$
and obtain a semigroup of linear operators $\{P^{X}_{t}\}_{t\geq 0}$ on $C_{\mathrm{Lip}}(\mathbb{R}^{N};\mathbb{R})$.
We begin by presenting an approximation theorem for the high-order discretisation scheme based on [11, 12, 13].
@@
Let $\ell$ be an integer. The vector fields $V_{0},V_{1},\dots,V_{d}$ are said to satisfy the UFG condition if for all $\alpha\in A^{\ast}_{1}$ there exists a family
$\left\{\varphi_{\alpha,\beta}\mid\beta\in A^{\ast}_{1}(\ell)\right\}\subset C^{\infty}_{b}(\mathbb{R}^{N};\mathbb{R})$
such that
@@
Hereafter, the integer $\ell$ in the definition above is referred to as the degree of the UFG condition.
We next introduce a theorem from [12], which concerns the smoothness of the semigroup $\{P^{X}_{t}\}_{t\geq 0}$ in certain directions.
Theorem 1.
Assume that the vector fields $V_{0},V_{1},\dots,V_{d}$ satisfy the UFG condition and let $X$ be the diffusion process defined in (1). Let $\alpha_{1},\alpha_{2},\dots,\alpha_{k}\in A^{\ast}_{1}$. Then there exists a constant $C>0$ such that
$\left\|V_{[\alpha_{1}]}\cdots V_{[\alpha_{k^{\prime}}]}P^{X}_{t}V_{[\alpha_{k^{\prime}+1}]}\cdots V_{[\alpha_{k}]}f\right\|_{L^{p}(dx)}\leq Ct^{\left(\|\alpha_{1}\|+\cdots+\|\alpha_{k}\|\right)/2}\|f\|_{L^{p}(dx)}.$
A condition on approximation operators is now introduced.
@@
A Markov operator $\tilde{Q}_{(s)}$ is said to be an $m$-similar operator if there exist ${\mathcal{L}}_{\mathbb{R}}(A)$-valued random variables $L_{(s)}^{(1)}$, $L_{(s)}^{(2)}$, $\dots$, and $L_{(s)}^{(K)}$ such that
@@
$E\left[{\left\|j_{m}\left(L_{(s)}^{(k)}\right)\right\|_{2}}^{n}\right]<\infty,\quad\left\langle L_{(s)}^{(k)},(0)\right\rangle=1\quad\text{for all $n\in\mathbb{N}$, $k\in\{1,2,\dots,K\}$},$
@@
The term $B^{\circ\alpha}(s)$ appearing above is defined for $\alpha\in A^{*}$ as follows:
$B^{\circ\alpha}(t)=\begin{cases}1&\text{if $\alpha=\emptyset$,}\\ \displaystyle{\int_{0}^{t}B^{\alpha^{\prime}}(s)\,\circ\mathrm{d}B^{a}(s)}&\text{if $\alpha=\alpha^{\prime}a$, $\alpha^{\prime}\in A^{*}$, $a\in A.$}\end{cases}$
The $m$-similar operator may be intuitively understood as an approximation of $E[f(X(T,x))]$ up to the $m$th order based on the stochastic Taylor expansion of the solution to the SDE (1). We refer to $m$ as the degree of the discretisation method. Note that this definition is a slight generalisation of the original definition given in [11, 12, 13]. The following theorem, due to Kusuoka [13], forms the foundation of high-order discretisation schemes.
Theorem 2.
Assume that the vector fields in the SDE (1) satisfy the UFG condition, that $X$ is the diffusion process defined by (1), and that $Q_{(s)}$ is an $m$-similar operator. Then there exist positive constants $C$ and $C^{\prime}$ depending only on $m$ and the vector fields $V_{0},V_{1},\dots,V_{d}$ such that
$\left\|P^{X}_{s}f-Q_{(s)}f\right\|_{\infty}\leq Cs^{(m+1)/2}\|f\|_{\mathrm{Lip}},$
@@
$\left\|Q_{(s)}f\right\|_{\infty}\leq\exp(C^{\prime}s)\|f\|_{\infty}$
for all $f\in C_{\mathrm{Lip}}(\mathbb{R}^{N};\mathbb{R})$.
Consequently, the total discretisation error is bounded as follows. Let $T>0$ be fixed, and consider a partition $\Delta$ of the interval $[0,T]$ with $\sharp\Delta=n$, given by $t_{0}=0<t_{1}<\dots<t_{n}=T$. Then there exists a constant $C^{\prime\prime}$, depending only on $m$, the vector fields $V_{0},V_{1},\dots,V_{d}$, and $T$ such that
$\left\|P^{X}_{T}f-Q_{(s_{n})}\circ Q_{(s_{n-1})}\circ\cdots\circ Q_{(s_{1})}f\right\|_{\infty}\\ \leq C^{\prime\prime}\|f\|_{\mathrm{Lip}}\left({s_{n}}^{1/2}+\sum_{j=m}^{m+1}\sum_{i=1}^{n-1}\frac{{s_{i}}^{(j+1)/2}}{(T-t_{i})^{j/2}}\right),$ (21)
@@
In the original work of Kusuoka [11, 13], Theorem 2 is proved under the assumption that $f\in C_{b}^{\infty}(\mathbb{R}^{N};\mathbb{R})$. This assumption can, however, be relaxed to $f\in C_{\mathrm{Lip}}(\mathbb{R}^{N};\mathbb{R})$, as pointed out in Section 3 of [18].